A wheel, 56 in. in diameter, is connected to a smaller one, 6 in. in diameter. They are connected to a spindle, on the end another wheel, 42 in. in diameter, is connected to another, 4 in. in diameter. Determine rpms of the latter wheel, the larger moving at 50 rpms.
My attempt:
Let x = rpms.
4/42 = 50/x
4x = 2100
x = 525 rpm.

42 in. = 3.5 ft.
4 in. = .33 ft.
.33/3.5 = 50/x
.33x = 175
x = 530.30 = 530 rpms.
Answer given: 4900
Where am I incorrect? Non-homework.





Answer by Alan3354(69443)

A wheel, 56 in. in diameter, is connected to a smaller one, 6 in. in diameter. They are connected to a spindle, on the end another wheel, 42 in. in diameter, is connected to another, 4 in. in diameter. Determine rpms of the latter wheel, the larger moving at 50 rpms.
My attempt:
Let x = rpms.
4/42 = 50/x
4x = 2100
x = 525 rpm.

42 in. = 3.5 ft.
4 in. = .33 ft.
.33/3.5 = 50/x
.33x = 175
x = 530.30 = 530 rpms.
Answer given: 4900
Where am I incorrect? Non-homework.
===================================
The arrangement is not clear, not to me at least.
Can you send a sketch? Or at least label the wheels A, B, C, etc., and not which are on the same shaft, if any?
------------
Send to gsihoutx@aol.com